OpenStudy (anonymous):

How would I multiply the following rational expression? ((a-b) (a+b))/((a^2-b^2))/((b)/(a+b))

OpenStudy (anonymous):

focus on each individual variable, so look at the numerator first: (a-b)(a+b). focus on the first parentheses and multiply each term into the next parentheses. a* a = a^2 a*b = ab then focus on the second term in the first parentheses -b*a = -ab -b*b = -b^2 combine all of them a^2 + ab - ab - b^2 the ab's cancel out a^2 -b^2 on the numerator that crosses out the dividing term under that to become one over b over a+b...i think. can you draw out your question?

OpenStudy (anonymous):